OpenStudy (anonymous):

use implicit differentiation to find y' if lnxy +5x=30

OpenStudy (amistre64):

use it then ... what part are you getting stuck at?

OpenStudy (anonymous):

the lnxy

OpenStudy (amistre64):

consider lnu, how would you approach its derivative?

OpenStudy (anonymous):

1/u(u')

OpenStudy (amistre64):

good, but to avoid some amibuity: u'/u let u = xy what is u'? think product rule

OpenStudy (amistre64):

**ambiguity...

OpenStudy (primeralph):

Or you can say ln(xy) = ln(x) + ln(y).

OpenStudy (anonymous):

ok so 1/x + 1/y(y')

OpenStudy (primeralph):

y'/y. Yes.

OpenStudy (amistre64):

my way: u = xy u' = x'y + xy' = y + xy' but yes, if you go the lnx + lny approach that would be fine as well

OpenStudy (anonymous):

and5x is just ln 1/x correct?

OpenStudy (primeralph):

No.

OpenStudy (anonymous):

or 5

OpenStudy (primeralph):

Of course.

OpenStudy (anonymous):

30 is 0

OpenStudy (anonymous):

so i got -y-5xy/x